### v2

# Open-Source ESP32-C3 Wi-Fi Drivers for Static Analysis

Abstract.
The Battery-Free Internet of Things will revolutionize our understanding of sustainable communication, as it operates on energy harvested from the environment (e.g., by solar cells). On these systems, harvested energy is dependent on unpredictable environmental conditions; therefore, device operations, including those of its networking stack, must be resilient to power failures. Reactive intermittent computing provides an approach for solving this by notifications of impending power failures, which is implemented by monitoring the harvested energy buffered in a capacitor. However, to use this power-failure notification and guarantee forward progress, systems must break down tasks into atomic sub-tasks (transactions) that can be predictably finished before the available energy runs out. Thus, static program-code analysis must determine the worst-case energy consumption (WCEC) of all transactions. In the case of Wi-Fi–capable devices, drivers are often closed-source, and it is therefore unclear whether their transactions have a WCEC since static analysis requires all code and its semantics.
In this work, we integrate a transactional networking stack with reverse-engineered Wi-Fi drivers to enable full-stack WCEC analysis for physical transmission and reception of packets. Further, we extended a static worst-case analysis approach with a resource-consumption model of our Wi-Fi driver. Our evaluation with the RISC-V–based ESP32-C3 platform gives worst-case bounds with our static analysis for the transactions of the full communications stack, therefore showing that Wi-Fi–based reactive intermittent computing is feasible.
1. Introduction
Battery-Free Systems
Battery-free devices (Ahmed et al., 2024) rely on energy harvested from various sources, such as light, radio-frequency signals, or vibrations. Being battery-free not only results in independence from battery requirements but also makes novel energy-self–sufficient applications possible: Battery-free medical implants (Kim et al., 2024) could reduce the need for surgical replacements due to degradation in battery performance, which puts such scenarios in the realm of both safety- and security-critical systems (Mottola et al., 2024). Naturally, the unpredictable nature of harvesting energy means that a started task can face a power failure before reaching a checkpoint, stored in non-volatile memory. These power failures cause two major problems: First, such a failure can leave the system in an inconsistent state. Second, when power failures occur during energy-intensive operations with transactional semantics (e.g., sending out a packet over a transceiver), the system might repeatedly face power failures, leaving the system in a livelock-like situation.
Worst-Case Analysis
To solve these two problems of inconsistency and starvation, the use of static worst-case program analysis can give guarantees on the execution of operations. Specifically, static worst-case energy consumption (WCEC) analysis techniques (Choi et al., 2022; Raffeck et al., 2024; Yarahmadi and Rohou, 2020) can determine a safe upper bound on the required energy demand of operations. These static program analysis techniques originally stem from the real-time systems community (Wilhelm et al., 2008) and their requirement to determine worst-case execution time (WCET) estimates for guaranteeing timeliness. With WCEC estimates and knowledge during runtime of the available energy in the storage (i.e., capacitor), operations run without the risk of power failures. Thus, consistency and forward progress is guaranteed.
Available Code & Semantics
A necessary requirement for applying such static resource-consumption analyses is the availability of all code along with its resource-consumption–related semantics (Abella et al., 2015; Schoeberl, 2009). This required knowledge is particularly important for power-consuming operations, such as radio transceivers, since inaccuracies in the static behavior modeling can result in (1) unsound or (2) impractically pessimistic analysis estimates. Unfortunately, for our embedded target platform, the RISC-V–based ESP32-C3 system-on-chip (SoC) from Espressif, the Wi-Fi drivers are closed-source, and the Wi-Fi module’s behavior is undocumented on the granularity that we require for static analysis.
Contributions
Fortunately, sophisticated tooling support from the security community enabled us to reverse engineer the necessary knowledge we require for conducting worst-case analyses. Further, the existing ESP32-Open-MAC project (Devreker, 2023), targeting Tensilica’s Xtensa SoC ESP32, allows us to benefit from an existing infrastructure. In summary, this work’s contributions are threefold:
Case Study on Reverse Engineering: We present our development of a Wi-Fi stack for the ESP32-C3, relying on static reverse-engineering analysis techniques (i.e., Ghidra) combined with dynamic analysis (i.e., binary-patched executables). The stack is designed to be fully statically analyzable and therefore suited for intermittently-powered devices.
Energy-Consumption & Execution-Time Model for Worst-Case Analysis: Our static analysis comes with a fine-grained energy-consumption model of the Wi-Fi driver for the ESP32-C3 target platform.
Open-Source Implementation & Evaluation: Based on our publicly available driver for ESP32-C3, we present evaluation results that validate the stack’s property of being statically analyzable.
As part of this work, we envision that concerns from the security domain, such as security auditability of the Wi-Fi stack’s behavior, can be interconnected with questions from the static-analysis domain, such as the resource demand of the Wi-Fi stack’s behavior.
Paper’s Organization
Section 2 introduces the necessary background information for our domain of Wi-Fi–enabled battery-free systems. Our main problems are stated in Section 3, while Section 4 gives insight into how we solved these problems. Section 5 presents our evaluations on a widely used embedded platform (i.e., ESP32-C3). Related work is discussed in Section 6. Section 7 outlines future work and concludes the paper.
2. Background
The following section first gives some background regarding embedded network stacks (see Section 2.1). In section 2.2, we discuss mechanisms through which processors and Wi-Fi peripherals communicate in low-power SoCs. Finally, 2.3 describes the approach of worst-case resource-consumption analysis and its role in intermittent systems.
2.1. Network stack
The network stack on embedded devices typically consists of a lightweight TCP/IP implementation and a driver that interacts with the physical network interface (e.g., Wi-Fi peripheral). At the highest level of the stack is the application that provides data to be transmitted, while at the lowest level is the physical layer (PHY), which sends out the data onto the physical medium. On top of the physical layer is the medium access control (MAC) that coordinates transmission between multiple devices. The MAC layer is tasked with network association, collision avoidance, or data encapsulation. In the case of Wi-Fi, collision avoidance (CSMA/CA) mechanism is used to prevent multiple devices from transmitting at the same time. Each layer of the network stack encapsulates data from the upper layer based on network configuration and passes it to the layer below. On intermittent devices, power may be lost at any point before packet transmission is completed, leading to loss of progress up to that point.
In the case of Wi-Fi, PHY operations are implemented in hardware due to strict timing requirements. Wi-Fi peripherals that use the FullMAC architecture may implement both MAC and PHY layer operations in hardware. In some peripherals, tasks such as network association may be implemented in software (SoftMAC).
Popularly, Wi-Fi devices operate in the basic service set (BSS) topology and connect to an access point (AP) that arbitrates communication within the network. Once connected, the devices are known as stations.
2.2. Wi-Fi Hardware
High-speed communication between the Wi-Fi peripheral and the processor within an SoC is typically carried out using a combination of memory-mapped I/O (MMIO) and direct memory access (DMA) (Schulz, 2018). The Wi-Fi peripheral is configured through MMIO, while data transfer for transmission and reception is performed using DMA. DMA allows the CPU to remain free for other tasks instead of handling the transfer. Espressif devices such as ESP32-C3 follow this architecture. The peripheral communicates events such as successful transmission or reception of a packet through interrupts. On the other hand, external Wi-Fi chipsets may use the SPI interface for communication with the host controller. Besides the ESP32-C3’s Wi-Fi chip, the CYW43439 is another example suitable for low-power systems (Cypress Semiconductor Corporation, 2021).
2.3. Static Worst-Case Analysis
Static worst-case resource-consumption analysis comprises two major pillars: (1) a hardware-agnostic path analysis and (2) a hardware-dependent cost modeling capturing the actual resource demand (i.e., execution time or energy demand) on the target platform. The main idea is to formulate a mathematical problem formulation, usually an integer linear program (ILP): The ILP’s constraints originate from the path analysis. The ILP’s objective function states the maximization of the actual costs with the given path constraints. Eventually, the ILP’s solution yields a mathematically proven worst-case bound, which, in turn, is necessary to ensure runtime guarantees.
3. Problem Statement
The problems that we address in this work are twofold: Section 3.1 discusses the problem of closed-source drivers, and Section 3.2 explains the requirement of resource models.
3.1. Problem of Closed-Source Wi-Fi Drivers
Our approach to intermittently-powered devices exploits static worst-case analysis techniques to gain knowledge on the worst-case energy consumption of operations. During runtime, our system’s scheduler checks the available energy and dispatches tasks only if sufficient energy is available to guarantee freedom from power failures. This WCEC-aware approach gives both memory-consistency and forward-progress guarantees. The trade-off we exploit is favoring expensive design-time analyses over complicated runtime checkpointing mechanisms. Once deployed, systems achieve provable resilience against power failures. However, these design-time analyses require the code to be available. Our target platform is the ESP32-C3, a RISC-V–based, widely used, and low-cost SoC. Although BLE, likewise available on the ESP32-C3, is less power-intensive compared to Wi-Fi, we are interested in the more challenging Wi-Fi stack being of high relevance: In 2024, an estimated 31 % of IoT devices used Wi-Fi for connectivity (Sinha, 2024). For this ESP32-C3 platform, the Wi-Fi drivers were not available as analyzable source code, which is a common problem for embedded SoC platforms. Of particular relevance for our analysis are the MAC and PHY layer operations.
Our Approach: To solve the problem of the closed source, our approach exploits several reverse-engineering techniques on Espressif’s closed-source Wi-Fi driver. Fortunately, Espressif legally allows reverse engineering this driver. Based on the reverse-engineering results, we contribute an open-source driver for the ESP32-C3.
3.2. Missing Resource-Consumption Model
Although the developed open-source driver gives us the freedom to perform static analysis,
a major problem is connecting memory-mapped IO operations to semantics with regard to the power demand.
When considering the following code for modifying memory
*((uint32_t*)0x60033ca0) &= 0xff00efff
the static analysis is not aware of the semantics of this memory access.
On the ESP32-C3 target platform, this operation switches on the transceiver from a sleep state to standby mode (see also Figure 1).
As a consequence thereof, our static analysis of this memory access requires the notion that the power drawn increases after this operation.
Similarly, specific instructions also require a longer execution time (than a normal RAM access), which in turn leads to higher energy demand (i.e., power over time).
Consequently, we require a model that comprehensively accounts for power, energy, and time.
Our Approach: For our target, we present a resource-consumption model that considers transitions in the hardware that change the power demand of the entire SoC. Our analysis makes use of this model and tracks program paths in a context-sensitive way.
4. Approach
We use a combination of commonly available static and dynamic reverse-engineering analysis methods for identifying the registers and correspondingly their contents. Espressif provides their SoftMAC implementation as binaries under the permissive Apache license. The binaries also contain the symbol table.
4.1. Reverse Engineering
Static Analysis
Through decompilation, the execution flow and accordingly the registers used for transmission and reception can be identified. Functions within the binaries are named based on their purpose, which leads to informative decompilation results. Additionally, the interrupt configuration used by the Wi-Fi peripheral can be identified as this is not exposed to developers.
However, decompilation alone is insufficient to identify control and data flows. Several of the registers are configured based on data that is internal to the driver, whose structure is not documented. Moreover, the peripheral also has several slots for transmission, which is also dynamically selected based on the network configuration. Therefore, we use QEMU to dynamically determine the register contents as well as the slots.
Dynamic Analysis
The QEMU version in (lcgamboa, 2023), a fork of Espressif’s QEMU, supports Wi-Fi for ESP32-C3 by identifying a select set of registers and emulating an open Wi-Fi AP. We modify this fork to log all peripheral accesses and their contents. This allows us to verify register addresses as well as determine the register operations corresponding to the Wi-Fi network configuration in QEMU.
While QEMU provides insights into all the peripheral’s register accesses and their contents, its emulation does not fully account for peripheral configuration. It only emulates the DMA transfer as well as the interrupt mechanism. Moreover, it is not always possible to deduce the purpose of the register operation.
Binary Patching for Injecting Tracing Code
We modify the control flow to trace execution in real-time. We make no modifications to the binaries themselves but replace exported symbol names within some binary files. The new symbol name can then be declared as a function as part of a regular ESP-IDF sketch and can call the original function with the provided arguments. This also allows us to peek into the arguments themselves without using breakpoints or external debuggers. The resulting binary can also be run on the emulator or on the ESP32-C3. However, this only allows us to create such function redirects only for those functions that are declared in one object file but called from another.
These reverse-engineering methods in combination with the existing infrastructure in ESP32-Open-MAC yield necessary information for constructing the Wi-Fi driver for ESP32-C3.
[device graph]device graph showing energy states
4.2. The Open-Source ESP32-C3 Driver
Once the registers and transmit/receive mechanisms are identified, we use the software architecture shown in Figure 2 to develop the driver.
MAC Task
The MAC task accepts packets in Ethernet (802.3) format from the IP layer. It adds the Wi-Fi MAC header and passes it to the driver for transmission. On the other hand, packets arriving from the driver are passed to the IP layer through a callback (CB). The MAC task provides a transmit primitive for the IP stack to queue packets for transmission. Conversely, a receive callback has to be provided by the IP stack.
Driver Task
The driver task interacts with the MMIO registers. It also manages the memory allocated to DMA structures required by the peripheral. All memory is statically allocated to aid WCET/C analysis. Processing of interrupts is deferred by the interrupt handler to the driver task. Incoming 802.11 packets are passed to the MAC layer for further processing. The driver task also defines two queues: a hardware event queue and a MAC event queue. The interrupt handler notifies the driver of the occurrence and reason for the interrupt using the hardware event queue. The MAC task also writes packets to be transmitted into this queue. On the other hand, the driver task writes any received packets into the MAC event queue to signal the MAC task of received packets.
[architecture]architecture of firmware
4.3. Building a Resource-Consumption Model
For tasks such as transmission and reception, the Wi-Fi peripheral functions independently of the CPU. Therefore, static code analysis has to be supplemented with resource consumption models, namely time and energy.
Timing Information
The Wi-Fi peripheral performs packet transmission and notifies the CPU of the status through interrupts. Similarly, some MMIO register operations take time due to ongoing receive tasks. However, these tasks are still completed or fail within a finite amount of time and therefore have upper bounds on their execution times.
Energy Information
The device energy model should consist of all possible energy states, as well as the corresponding energy consumption. In our current setup, we assume that the CPU is in the always-on state; however, modeling the ESP32-C3’s sleep states, (low-power) run states, and frequency changes is possible (Dengler et al., 2023). For the Wi-Fi peripheral, we define three states of operation: Sleep, Standby and Transmitting. The Wi-Fi peripheral is powered-down in and therefore consumes no energy. In , the Wi-Fi peripheral is powered up and may receive packets, thereby consuming energy. In , the device transmits packets. The energy consumption varies based on the duration of transmission. No reception occurs in this state, as the peripheral owns the channel through the CSMA-CA mechanism. The device state machine is shown in Figure 1.
4.4. Implementation
Reverse Engineering
The ESP32-Open-MAC project lists the registers that are necessary to achieve basic transmission and reception on the ESP32. In order to identify the equivalent registers on the ESP32-C3, we build a sketch that performs simple UDP transmission and analyze it using Ghidra’s (Agency, 2019) decompiler. We also add volatile memory sections for the peripheral ranges identified in previous reverse engineering efforts (lcgamboa, 2023). The author also provides a fork of Espressif’s QEMU that adds support for Wi-Fi functionality on the ESP32-C3. We modify this fork to log all peripheral accesses.
We replace only the external function names in the binaries with function names of similar length. Furthermore, we declare these functions in the main ESP-IDF sketch to the original function names in the object files. The resulting sketch is both loaded to the emulator and flashed to an ESP32-C3-Devkit-M1.
| Register | Deduced purpose | Access |
| 0x60033000 | MAC address configuration | R/W |
| 0x60033084 | MS bit enables or disables Rx | R/W |
| 0x60033088 | Set base DMA linked list address | R/W |
| 0x60033ca0 | Power up and power down Wi-Fi module | R/W |
| 0x60033d08 | Configure Tx DMA address and trigger Tx | R/W |
| 0x60033c3c | Get interrupt reason | R |
| 0x60033c40 | Clear interrupt | R/W |
| 0x60033cac | Clear Tx slot | R/W |
Some important registers and their deduced purpose are shown in Table 1.
Interrupt Handling
Setting up interrupts on the ESP32-C3 is well documented. However, the Wi-Fi interrupt is not exposed to developers. Through decompilation, we find that the Wi-Fi driver uses CPU interrupt number 1. To register a custom Wi-Fi interrupt handler, we find that first the source register corresponding to interrupt 1 (0x600c2000) has to be cleared. The interrupt has to be disabled by masking bit 1 of 0x600c2104. The custom interrupt handler can be set using intr_handler_set provided by ESP-IDF. Finally, the interrupt is enabled again by setting bit 1 of 0x600c2104.
ESP32-C3 Wi-Fi driver
We implement the driver using the architecture described in Figure 2. We use an open AP, as the encryption functionality of Wi-Fi peripheral hardware is not fully reverse-engineered. For initialization and connecting to an AP, we use the API calls esp_wifi_init and esp_wifi_start provided by ESP-IDF.
These calls initialize and configure the Wi-Fi peripheral, set the internal interrupt handler for Wi-Fi, and initiate a connection to the AP. The successful completion of these tasks can be monitored through a system notification WIFI_EVENT_STA_CONNECTED. Once the notification is received, we kill the Wi-Fi task using the undocumented pp_post function and set up the custom interrupt handler. Finally, our driver and MAC tasks are started. We use PfIP (Vogelgesang et al., 2025) as the IP stack due to its intermittency-aware execution.
Static Analysis
For static analysis, we use the Platin framework (Maroun et al., 2024). Platin uses a modified version of the clang compiler that collects program flow information. The compiler generates a metadata file from the program flow information and transforms it into an ILP, which, in turn, can be solved using programs such as lp_solve (Berkelaar et al., 2004) or gurobi (Gurobi Optimization, LLC, 2023). For the purpose of this paper, we annotate the current consumption for each state in the device’s energy model based on the datasheet (Espressif Systems, 2022). Additionally, Platin distinguishes between the energy penalty of switching on and off any given energy state (wcec-turn-on and wcec-turn-off) and the energy consumption of a device in the state energy-stay-on. wcec-always-on represents the energy required when the device is in any given state. In our model, we set both the switching on and off penalties to zero and model the device-switching behavior pessimistically.
5. Evaluation
This section presents the evaluation of our open-source Wi-Fi driver. Section 5.1 describes the hardware and network setup as well as the task system used for the experiments. Additionally, we detail the required assumptions for the static resource-consumption analysis using the Platin analysis framework (Maroun et al., 2024). In Section 5.2, we then examine and discuss the results from the static analysis.
5.1. Evaluation Setup
Hardware Setup
[…]
Functionality
To validate the functionality of the Wi-Fi driver, we perform an experiment with three tasks. The mac_task interacts with the PfIP network stack, the hw_task initiates the transmission of packets required by the mac_task and handles receiving, and a test_task sends out a packet using the PfIP stack and waits for an answer from other devices in the network. We validated the functionality of our open-source driver with Wireshark captures of the network traffic.
Static Analysis
To analyze the worst-case resource demand of our Wi-Fi driver, we use the Platin analysis framework (Maroun et al., 2024) in combination with the device-model analysis of the WoCA approach (Raffeck et al., 2024). The focus of our evaluation lies solely on the driver functions, independent of the runtime and operating system. While a whole-system analysis, which includes the operating system, is generally possible (Schuster et al., 2019), we exclude the FreeRTOS components and the interaction with the FreeRTOS scheduler from our analysis. Likewise, we exclude the callbacks into the PfIP network stack from the analysis.
We perform the worst-case resource-demand analysis both for individual driver functions and for a task comprising multiple device interactions. For each individual driver function, we determine an upper bound on the execution time. The evaluation task transmits a packet and actively waits for the completion of the transmission. After starting the transmission, the Wi-Fi transceiver transmits the packet and waits for an acknowledgment (ACK). Once the transmission is completed, the transmission returns to standby mode. We derive an upper bound on both the time and energy demand with respect to the different device states of the Wi-Fi transceiver during the task execution.
[…]
5.2. Evaluation Results
| WCET | always-on WCEC | ||
| wifi_hw_deinit | 48 cy | 0.3 us | 0.3099 uJ |
| wifi_setup_interrupt | 178 cy | 1.11 us | 1.149 uJ |
| wifi_setup_rx | 1881 cy | 11.8 us | 12.14 uJ |
| wifi_hw_init | 49 cy | 0.306 us | 0.3163 uJ |
| wifi_transmit_packet | 335 cy | 2.09 us | 2.163 uJ |
| wifi_wait_for_tx | 52184 cy | 326 us | 336.9 uJ |
| wifi_process_tx_done | 157 cy | 0.981 us | 1.014 uJ |
| wifi_handle_rx | 12989 cy | 81.2 us | 83.85 uJ |
| wifi_process_timeout | 138 cy | 0.862 us | 0.8909 uJ |
| wifi_get_bssid | 94 cy | 0.588 us | 0.6068 uJ |
| wifi_mac_handle_rx | 68715 cy | 429 us | 443.6 uJ |
| wifi_interrupt_handler | 943 cy | 5.89 us | 6.088 uJ |
| WCET | always-on WCEC | device-aware WCEC | ||
| TX task | 52615 cy | 329 us | 339.7 uJ | 32.42 uJ |
[…]
Most functions of our open-source Wi-Fi driver perform only operations on memory-mapped registers to control the transceiver and exhibit relatively low execution time and energy consumption. wifi_mac_handle_rx copies the payload data of the received package between buffers, so the analysis has to assume the worst case of the maximum packet size. Similarly, the analysis has to assume the worst-case number of packets to handle for the execution time of wifi_handle_rx and the maximum time until an ACK is received for the execution time of wifi_wait_for_tx. Finally, the difference between the always-on and device-aware WCEC bounds underlines the importance of considering the states of the Wi-Fi transceiver in a fine-granular manner.
6. Related Work
Typical approaches for battery-free wireless communication use Bluetooth Low Energy (BLE), LoRa , or backscatter-based wireless communications (Cai et al., 2023; Talla et al., 2017; Jiang et al., 2023). Intermittency-aware operation has been successfully demonstrated, especially for BLE and LoRa.
Intermittency-Aware Wireless Communication
Wireless networking is a key challenge in Battery-free IoT designs due to intermittency. Therefore, the networking stack deployed on these devices must be capable of intermittent execution. (Fu et al., 2023) proposed a fully intermittency-aware networking stack. Schoeberl et al. developed tpIP, a time-predictable IP stack for cyber-physical networks (Schoeberl and Pedersen, 2018). tpIP is designed to allow static analysis through bounded loops and avoids dynamic memory allocation. They also estimated WCET for the stack’s operations. Similarly, Vogelgesang et al. designed PfIP, a transactional networking stack with worst-case bounds for every transaction (Vogelgesang et al., 2025). These approaches are agnostic to the underlying physical layer.
De Winkel et al. introduced a battery-free smartwatch running FreeBie (de Winkel et al., 2022), a bidirectional BLE stack for intermittently-powered Bluetooth devices. The device stored checkpoints for all processes when power failure was imminent and restored them when power was available. Additionally, due to intermittency, two devices may not be awake at the same time for communication. In (Geissdoerfer and Zimmerling, 2021, 2022), these challenges are addressed through probabilistic modeling of latency and charge times, respectively.
Similarly, a LoRa-based battery-free image sensing system that adaptively configures its transmission mode based on the available energy was developed in (Nardello et al., 2019). Raffeck et al. demonstrated a LoRa-based device using WoCA, a whole system worst-case analysis approach for intermittent wireless devices (Raffeck et al., 2024). For static analysis, Platin was used with loop bounds calculated based on the LoRa device configuration.
Greentooth is a proposed low-power wireless protocol designed for battery-free devices (Babatunde et al., 2024). Greentooth synchronizes communication between two devices using time division multiple access (TDMA)-style scheduling and low-power wake-up radios.
Wi-Fi remains under-explored due to its higher power requirement compared to BLE or LoRa. The ubiquity of Wi-Fi networking applications and availability of off-the-shelf inexpensive hardware such as the ESP32-C3 promises battery-free innovation in the field of, for example, smart-home applications, where an existing Wi-Fi network is usually available.
Open-Source Wi-Fi Platforms
Open-source drivers exist for several Wi-Fi network interface cards from, for example, Atheros or Mediatek (Project, 2024). These require no binary blobs and are often supported by the manufacturers themselves. However, they require a Linux kernel and PCI-E hardware support and are generally not intended for low-power SoCs.
The OpenWiFi project (Jiao et al., 2020) developed an open-source FPGA-based software-defined radio for the 802.11 standard paired with an RF frontend. However, FPGAs are not commonly used in consumer hardware and the platform is intended to be a tool for researchers.
On low-power devices, the ZeroWi project (Bentham, 2020) focuses on reverse-engineering the capabilities of the Cypress CYW43438 Wi-Fi SoC. The CYW43438 consists of an internal ARM Cortex CPU, which is programmed using non-free firmware provided by Cypress. Drivers written for this chip then send commands in the SDIO format over SPI to communicate with the firmware.
For the Espressif products, early attempts at reverse engineering were performed by Uri Shaked, who used it as part of his Wokwi simulator for ESP32 (Shaked, 2019). The QEMU-based open-source PICSimLab simulator introduces support for Wi-Fi emulation on the ESP32 and ESP32-C3 devices (lcgamboa, 2022). Based on these attempts, the ESP32-Open-MAC project (Devreker, 2023) aims to fully open-source the Wi-Fi driver on the ESP32. Our work not only ported ESP32-Open-MAC onto the RISC-V–based ESP32-C3 but also integrated it with (1) a resource-consumption model for static analysis and (2) a transactional network stack that is suitable for battery-free devices.
7. Future Work & Conclusion
The current energy model does not consider packet reception. Therefore, future work needs to consider a more detailed device model: That is, pessimism in the bounds can be reduced by a more fine-grained device model. For example, in , the peripheral is not actively transmitting after the end of transmission until ACK is received. Additionally, energy penalties for switching on/off energy states could be accurately considered. Further, we use a static peripheral configuration. The resource-consumption bounds are impacted by peripheral configurations, such as data rate of transmission. An interesting avenue for future work is to identify optimal configurations under energy constraints. The ESP32-Open-MAC project also has recently introduced initialization independent of FreeRTOS. Therefore, an OS-aware implementation could also result in easier static analysis, for example, by exploiting Rust’s statically analyzable async/await pattern.
In this work, we presented a fully open-source Wi-Fi driver for the ESP32-C3, based on the ESP32-Open-MAC project. We successfully port the project from the Tensilica-based ESP32 to the RISC-V–based ESP32-C3 using common reverse-engineering techniques. Finally, we present a proof-of-concept for performing static resource-consumption analysis on the physical operations of the Wi-Fi peripheral using the extended Platin worst-case analysis toolkit. For the future, we envision more exchange between the static-analysis and the security communities. More interdisciplinary research on this topic could avoid a future of insecure and unsafe battery-free IoT systems.
[ backgroundcolor=green!10, linecolor=green!70,linewidth=2pt,topline=false,bottomline=false, innertopmargin=0pt, innerbottommargin=5pt, innerleftmargin=6pt, innerrightmargin=6pt, frametitleaboveskip=5pt, frametitlefont=, frametitlealignment=, frametitle=The implementation is available at: https://git.nt.uni-saarland.de/open-access/pfip ]
Acknowledgements.
This work is supported by the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG) as part of SPP 2378 (Resilient Worlds: project number 502615015, Resilient Power-Constrained Embedded Communication Terminals, ResPECT). We would also like to thank all ESP32-Open-MAC contributors for their valuable contributions to the community.References
- (1)
- Abella et al. (2015) Jaume Abella, Carles Hernández, Eduardo Quiñones, Francisco J. Cazorla, Philippa Ryan Conmy, Mikel Azkarate-askasua, Jon Pérez, Enrico Mezzetti, and Tullio Vardanega. 2015. WCET analysis methods: Pitfalls and challenges on their trustworthiness. In Proceedings of the 10th International Symposium on Industrial Embedded Systems (SIES ’15). IEEE, 39–48. https://doi.org/10.1109/SIES.2015.7185039
- Agency (2019) National Security Agency. 2019. Ghidra. https://ghidra-sre.org/.
- Ahmed et al. (2024) Saad Ahmed, Bashima Islam, Kasim Sinan Yildirim, Marco Zimmerling, Przemyslaw Pawelczak, Muhammad Hamad Alizai, Brandon Lucia, Luca Mottola, Jacob Sorber, and Josiah D. Hester. 2024. The Internet of Batteryless Things. Commun. ACM 67, 3 (2024), 64–73. https://doi.org/10.1145/3624718
- Babatunde et al. (2024) Simeon Babatunde, Arwa Alsubhi, Josiah D. Hester, and Jacob Sorber. 2024. Greentooth: Robust and Energy Efficient Wireless Networking for Batteryless Devices. ACM Transactions on Sensor Networks 20, 3 (may 2024), 66:1–66:31. https://doi.org/10.1145/3649221
- Bentham (2020) Jeremy Bentham. 2020. Zerowi: Bare-Metal WiFi Driver for the Raspberry Pi. https://iosoft.blog/zerowi/
- Berkelaar et al. (2004) Michel Berkelaar, Kjell Eikland, and Peter Notebaert. 2004. lp_solve 5.5, open source (mixed-integer) linear programming system. Software.
- Cai et al. (2023) Zhipeng Cai, Quan Chen, Tuo Shi, Tongxin Zhu, Kunyi Chen, and Yingshu Li. 2023. Battery-Free Wireless Sensor Networks: A Comprehensive Survey. IEEE Internet of Things Journal 10, 6, March 15 (mar 2023), 5543–5570. https://doi.org/10.1109/JIOT.2022.3222386
- Choi et al. (2022) Jongouk Choi, Larry Kittinger, Qingrui Liu, and Changhee Jung. 2022. Compiler-Directed High-Performance Intermittent Computation with Power Failure Immunity. In Proceedings of the 28th Real-Time and Embedded Technology and Applications Symposium (RTAS ’22). IEEE, 40–54. https://doi.org/10.1109/RTAS54340.2022.00012
- Cypress Semiconductor Corporation (2021) Cypress Semiconductor Corporation 2021. CYW43439 Single-Chip IEEE 802.11 b/g/n MAC/Baseband/Radio with Integrated Bluetooth 5.2 Compliance. Cypress Semiconductor Corporation. https://www.mouser.com/datasheet/2/196/Infineon_CYW43439_DataSheet_v03_00_EN-3074791.pdf
- de Winkel et al. (2022) Jasper de Winkel, Haozhe Tang, and Przemyslaw Pawelczak. 2022. Intermittently-Powered Bluetooth That Works. In Proceedings of the 20th Annual International Conference on Mobile Systems, Applications and Services (MobiSys ’22) (Portland, Oregon) (MobiSys ’22). ACM, New York, NY, USA, 287–301. https://doi.org/10.1145/3498361.3538934
- Dengler et al. (2023) Eva Dengler, Phillip Raffeck, Simon Schuster, and Peter Wägemann. 2023. FusionClock: Energy-Optimal Clock-Tree Reconfigurations for Energy-Constrained Real-Time Systems. In Proceedings of the 35th Euromicro Conference on Real-Time Systems (ECRTS ’23), Vol. 262. 6:1–6:24. https://drops.dagstuhl.de/opus/volltexte/2023/18035/pdf/LIPIcs-ECRTS-2023-6.pdf
- Devreker (2023) Jasper Devreker. 2023. Unveiling secrets of the ESP32: creating an open-source MAC Layer. Retrieved 2025-01-03 from https://zeus.ugent.be/blog/23-24/open-source-esp32-wifi-mac/
- Espressif Systems (2022) Espressif Systems 2022. ESP32-C3 Series Datasheet. Espressif Systems. https://www.espressif.com/sites/default/files/documentation/esp32-c3_datasheet_en.pdf
- Fu et al. (2023) Shen Fu, Vishak Narayanan, Mathew L. Wymore, Vishal Deep, Henry Duwe, and Daji Qiao. 2023. No battery, no problem: Challenges and opportunities in batteryless intermittent networks. Journal of Communications and Networks 25, 6 (2023), 806–813. https://doi.org/10.23919/jcn.2023.000033
- Geissdoerfer and Zimmerling (2021) Kai Geissdoerfer and Marco Zimmerling. 2021. Bootstrapping Battery-free Wireless Networks: Efficient Neighbor Discovery and Synchronization in the Face of Intermittency. In 18th USENIX Symposium on Networked Systems Design and Implementation (NSDI 21). USENIX Association, 439–455. https://www.usenix.org/conference/nsdi21/presentation/geissdoerfer
- Geissdoerfer and Zimmerling (2022) Kai Geissdoerfer and Marco Zimmerling. 2022. Learning to Communicate Effectively Between Battery-free Devices. In 19th USENIX Symposium on Networked Systems Design and Implementation (NSDI 22). USENIX Association, 419–435. https://www.usenix.org/conference/nsdi22/presentation/geissdoerfer
- Gurobi Optimization, LLC (2023) Gurobi Optimization, LLC. 2023. Reference Manual. gurobi.com/.
- Jiang et al. (2023) Tao Jiang, Yu Zhang, Wenyuan Ma, Miaoran Peng, Yuxiang Peng, Mingjie Feng, and Guanghua Liu. 2023. Backscatter Communication Meets Practical Battery-Free Internet of Things: A Survey and Outlook. IEEE Communications Surveys & Tutorials 25, 3 (2023), 2021–2051. https://doi.org/10.1109/COMST.2023.3278239
- Jiao et al. (2020) Xianjun Jiao, Wei Liu, Michael T. Mehari, Muhammad Aslam, and Ingrid Moerman. 2020. openwifi: a free and open-source IEEE802.11 SDR implementation on SoC. In 91st IEEE Vehicular Technology Conference, VTC Spring 2020, Antwerp, Belgium, May 25-28, 2020. IEEE, IEEE, 1–2. https://doi.org/10.1109/VTC2020-Spring48590.2020.9128614
- Kim et al. (2024) Young-Jun Kim, Jiho Lee, Joon-Ha Hwang, Youngwook Chung, Byung-Joon Park, Junho Kim, So-Hee Kim, Junseung Mun, Hong-Joon Yoon, Sung-Min Park, et al. 2024. High-Performing and Capacitive-Matched Triboelectric Implants Driven by Ultrasound. Advanced Materials 36, 2 (2024), 2307194. https://doi.org/10.1002/adma.202307194
- lcgamboa (2022) lcgamboa. 2022. Picsimlab. https://lcgamboa.github.io/
- lcgamboa (2023) lcgamboa. 2023. Qemu: Modified Version of the Espressif QEMU Used by the PICSimLab Simulator. https://github.com/lcgamboa/qemu/tree/picsimlab-esp32
- Maroun et al. (2024) Emad Jacob Maroun, Eva Dengler, Christian Dietrich, Stefan Hepp, Henriette Herzog, Benedikt Huber, Jens Knoop, Daniel Wiltsche-Prokesch, Peter P. Puschner, Phillip Raffeck, Martin Schoeberl, Simon Schuster, and Peter Wägemann. 2024. The Platin Multi-Target Worst-Case Analysis Tool. In Proceedings of the 22nd International Workshop on Worst-Case Execution Time Analysis (WCET ’24), Vol. 121. Schloss Dagstuhl - Leibniz-Zentrum für Informatik, 2:1–2:14. https://doi.org/10.4230/OASIcs.WCET.2024.2
- Mottola et al. (2024) Luca Mottola, Arslan Hameed, and Thiemo Voigt. 2024. Energy Attacks in the Battery-less Internet of Things: Directions for the Future. In Proceedings of the 17th European Workshop on Systems Security (EuroSec ’24). ACM, 29–36. https://doi.org/10.1145/3642974.3652283
- Nardello et al. (2019) Matteo Nardello, Harsh Desai, Davide Brunelli, and Brandon Lucia. 2019. Camaroptera: A batteryless long-range remote visual sensing system. In Proceedings of the 7th International Workshop on Energy Harvesting & Energy-Neutral Sensing Systems (EnsSys ’19). ACM, 8–14. https://doi.org/10.1145/3362053.3363491
- Project (2024) Linux Wireless Documentation Project. 2024. Existing Linux Wireless Drivers — Linux Wireless Documentation. https://wireless.docs.kernel.org/en/latest/en/users/drivers.html
- Raffeck et al. (2024) Phillip Raffeck, Johannes Maier, and Peter Wägemann. 2024. WoCA: Avoiding Intermittent Execution in Embedded Systems by Worst-Case Analyses with Device States. In Proceedings of the 25th ACM SIGPLAN/SIGBED International Conference on Languages, Compilers, and Tools for Embedded Systems (LCTES ’24). ACM, 83–94. https://doi.org/10.1145/3652032.3657569
- Schoeberl (2009) Martin Schoeberl. 2009. Time-predictable computer architecture. EURASIP Journal on Embedded Systems 2009 (2009), 1–17. https://doi.org/10.1155/2009/758480
- Schoeberl and Pedersen (2018) Martin Schoeberl and Rasmus Ulslev Pedersen. 2018. tpip: A time-predictable tcp/ip stack for cyber-physical systems. In 2018 IEEE 21st International Symposium on Real-Time Distributed Computing (ISORC). IEEE, IEEE Computer Society, 75–82. https://doi.org/10.1109/ISORC.2018.00018
- Schulz (2018) Matthias Schulz. 2018. Teaching Your Wireless Card New Tricks: Smartphone Performance and Security Enhancements Through Wi-Fi Firmware Modifications. Ph. D. Dissertation. Darmstadt University of Technology, Germany, Darmstadt. http://tuprints.ulb.tu-darmstadt.de/7243/
- Schuster et al. (2019) Simon Schuster, Peter Wägemann, Peter Ulbrich, and Wolfgang Schröder-Preikschat. 2019. Proving Real-Time Capability of Generic Operating Systems by System-Aware Timing Analysis. In Proceedings of the 25th Real-Time and Embedded Technology and Applications Symposium (RTAS ’19). IEEE, 318–330. https://doi.org/10.1109/RTAS.2019.00034
- Shaked (2019) Uri Shaked. 2019. Wokwi - World’s Most Advanced ESP32 Simulator. https://wokwi.com/
- Sinha (2024) Satyajit Sinha. 2024. State of IoT 2024: Number of Connected IoT Devices Growing 13% to 18.8 Billion Globally. https://iot-analytics.com/number-connected-iot-devices/.
- Systems (2016) Espressif Systems. 2016. ESP IoT Development Framework | Espressif Systems. https://www.espressif.com/en/products/sdks/esp-idf.
- Talla et al. (2017) Vamsi Talla, Bryce Kellogg, Shyamnath Gollakota, and Joshua R. Smith. 2017. Battery-Free Cellphone. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies 1, 2 (Jun 2017), 25:1–25:20. https://doi.org/10.1145/3090090
- Vogelgesang et al. (2025) Kai Vogelgesang, Ishwar Mudraje, Luis Gerhorst, Phillip Raffeck, Peter Wägemann, Thorsten Herfet, and Wolfgang Schröder-Preikschat. 2025. PfIP: A UDP/IP Transactional Network Stack for Power-Failure Resilience in Embedded Systems. In Proceedings of the Consumer Communications & Networking Conference (CCNC 2025).
- Wilhelm et al. (2008) Reinhard Wilhelm, Jakob Engblom, Andreas Ermedahl, Niklas Holsti, Stephan Thesing, David B. Whalley, Guillem Bernat, Christian Ferdinand, Reinhold Heckmann, Tulika Mitra, Frank Mueller, Isabelle Puaut, Peter P. Puschner, Jan Staschulat, and Per Stenström. 2008. The Worst-case Execution-time Problem – Overview of Methods and Survey of Tools. ACM Transactions on Embedded Computing Systems (ACM TECS) 7, 3 (2008), 36:1–36:53. https://doi.org/10.1145/1347375.1347389
- Wireshark Foundation (2025) Wireshark Foundation. 2025. free and open-source packet analyzer. https://www.wireshark.org/.
- Yarahmadi and Rohou (2020) Bahram Yarahmadi and Erven Rohou. 2020. Compiler optimizations for safe insertion of checkpoints in intermittently powered systems. In Proceedings of the 20th International Conference on Embedded Computer Systems: Architectures, Modeling, and Simulation (SAMOS ’20), Vol. 12471. Springer, 169–185. https://doi.org/10.1007/978-3-030-60939-9_12

### v3

@@
-Open-Source ESP32-C3 Wi-Fi Drivers for Static Analysis
+Reverse Engineering the ESP32-C3 Wi-Fi Drivers for Static Worst-Case Analysis of Intermittently-Powered Systems
 Abstract.
-The Battery-Free Internet of Things will revolutionize our understanding of sustainable communication, as it operates on energy harvested from the environment (e.g., by solar cells).
-On these systems, harvested energy is dependent on unpredictable environmental conditions; therefore, device operations, including those of its networking stack, must be resilient to power failures.
+The Internet of Batteryless Things revolutionizes sustainable communication as it operates on harvested energy.
+This harvested energy is dependent on unpredictable environmental conditions; therefore, device operations, including those of its networking stack, must be resilient to power failures.
 Reactive intermittent computing provides an approach for solving this by notifications of impending power failures, which is implemented by monitoring the harvested energy buffered in a capacitor.
-However, to use this power-failure notification and guarantee forward progress, systems must break down tasks into atomic sub-tasks (transactions) that can be predictably finished before the available energy runs out.
+However, to use this power-failure notification and guarantee forward progress, systems must break down tasks into atomic transactions that can be predictably finished before the energy runs out.
 Thus, static program-code analysis must determine the worst-case energy consumption (WCEC) of all transactions.
-In the case of Wi-Fi–capable devices, drivers are often closed-source, and it is therefore unclear whether their transactions have a WCEC since static analysis requires all code and its semantics.
-In this work, we integrate a transactional networking stack with reverse-engineered Wi-Fi drivers to enable full-stack WCEC analysis for physical transmission and reception of packets.
-Further, we extended a static worst-case analysis approach with a resource-consumption model of our Wi-Fi driver.
-Our evaluation with the RISC-V–based ESP32-C3 platform gives worst-case bounds with our static analysis for the transactions of the full communications stack, therefore showing that Wi-Fi–based reactive intermittent computing is feasible.
+In Wi-Fi–capable devices, drivers are often closed-source, which avoids the determination of WCEC bounds for transactions since static analysis requires all code along with its semantics.
+In this work, we integrate an energy-aware networking stack with reverse-engineered Wi-Fi drivers to enable full-stack WCEC analysis for physical transmission and reception of packets.
+Further, we extended a static worst-case analysis tool with a resource-consumption model of our Wi-Fi driver.
+Our evaluation with the RISC-V–based ESP32-C3 platform gives worst-case bounds with our static analysis approach for the transactions of the full communication stack, therefore showing that Wi-Fi–based reactive intermittent computing is feasible.
 1. Introduction
-Battery-Free Systems
-Battery-free devices (Ahmed et al., 2024) rely on energy harvested from various sources, such as light, radio-frequency signals, or vibrations.
-Being battery-free not only results in independence from battery requirements but also makes novel energy-self–sufficient applications possible: Battery-free medical implants (Kim et al., 2024) could reduce the need for surgical replacements due to degradation in battery performance, which puts such scenarios in the realm of both safety- and security-critical systems (Mottola et al., 2024).
-Naturally, the unpredictable nature of harvesting energy means that a started task can face a power failure before reaching a checkpoint, stored in non-volatile memory.
-These power failures cause two major problems: First, such a failure can leave the system in an inconsistent state.
-Second, when power failures occur during energy-intensive operations with transactional semantics (e.g., sending out a packet over a transceiver), the system might repeatedly face power failures, leaving the system in a livelock-like situation.
+Batteryless Systems
+Batteryless devices (Ahmed et al., 2024) rely on harvested energy.
+Being batteryless not only results in independence from battery requirements but also makes novel energy-self–sufficient applications possible: Batteryless medical implants (Kim et al., 2024) reduce the need for surgical replacements due to degradation in battery performance, putting them into the scope of safety- and security-critical systems (Mottola et al., 2024).
+Naturally, the unpredictable nature of harvesting energy means that a started task can face a power failure before reaching a checkpoint stored in non-volatile memory.
+Power failures cause two major problems: First, failures can leave the system in an inconsistent state.
+Second, with failures during energy-intensive operations with transactional semantics (e.g., sending packets), systems might repeatedly face failures without making progress.
 Worst-Case Analysis
-To solve these two problems of inconsistency and starvation, the use of static worst-case program analysis can give guarantees on the execution of operations.
-Specifically, static worst-case energy consumption (WCEC) analysis techniques (Choi et al., 2022; Raffeck et al., 2024; Yarahmadi and Rohou, 2020) can determine a safe upper bound on the required energy demand of operations.
-These static program analysis techniques originally stem from the real-time systems community (Wilhelm et al., 2008) and their requirement to determine worst-case execution time (WCET) estimates for guaranteeing timeliness.
-With WCEC estimates and knowledge during runtime of the available energy in the storage (i.e., capacitor), operations run without the risk of power failures.
-Thus, consistency and forward progress is guaranteed.
+To address inconsistency and starvation, static worst-case program analysis can guarantee power-failure–free execution in intermittent systems.
+Specifically, static worst-case energy consumption (WCEC) analysis (Choi et al., 2022; Raffeck et al., 2024; Wägemann et al., 2018; Yarahmadi and Rohou, 2020) determines a safe upper bound on operations’ energy demand.
+These analysis techniques originally stem from determining the worst-case execution time (WCET) to guarantee timeliness in real-time systems (Wilhelm et al., 2008).
+With WCEC estimates and the available energy in the storage (i.e., capacitor) during runtime, operations execute without the risk of power failures.
+Consequently, consistency and forward progress are guaranteed under consideration of the currently available energy.
 Available Code & Semantics
-A necessary requirement for applying such static resource-consumption analyses is the availability of all code along with its resource-consumption–related semantics (Abella et al., 2015; Schoeberl, 2009).
-This required knowledge is particularly important for power-consuming operations, such as radio transceivers, since inaccuracies in the static behavior modeling can result in (1) unsound or (2) impractically pessimistic analysis estimates.
-Unfortunately, for our embedded target platform, the RISC-V–based ESP32-C3 system-on-chip (SoC) from Espressif, the Wi-Fi drivers are closed-source, and the Wi-Fi module’s behavior is undocumented on the granularity that we require for static analysis.
+A necessary requirement for such static resource-consumption analyses is the availability of all code along with its resource-consumption semantics (Abella et al., 2015; Schoeberl, 2009).
+This knowledge is particularly important for power-consuming operations, such as radio transceivers, since inaccuracies in the static behavior modeling can result in unsound or impractically pessimistic analysis estimates.
+Unfortunately, for our target platform, the RISC-V–based ESP32-C3 system-on-chip (SoC), the Wi-Fi drivers are closed-source, and the Wi-Fi peripheral’s behavior is undocumented.
 Contributions
-Fortunately, sophisticated tooling support from the security community enabled us to reverse engineer the necessary knowledge we require for conducting worst-case analyses.
-Further, the existing ESP32-Open-MAC project (Devreker, 2023), targeting Tensilica’s Xtensa SoC ESP32, allows us to benefit from an existing infrastructure.
-In summary, this work’s contributions are threefold:
-Case Study on Reverse Engineering: We present our development of a Wi-Fi stack for the ESP32-C3, relying on static reverse-engineering analysis techniques (i.e., Ghidra) combined with dynamic analysis (i.e., binary-patched executables).
-The stack is designed to be fully statically analyzable and therefore suited for intermittently-powered devices.
-Energy-Consumption & Execution-Time Model for Worst-Case Analysis: Our static analysis comes with a fine-grained energy-consumption model of the Wi-Fi driver for the ESP32-C3 target platform.
-Open-Source Implementation & Evaluation: Based on our publicly available driver for ESP32-C3, we present evaluation results that validate the stack’s property of being statically analyzable.
-As part of this work, we envision that concerns from the security domain, such as security auditability of the Wi-Fi stack’s behavior, can be interconnected with questions from the static-analysis domain, such as the resource demand of the Wi-Fi stack’s behavior.
-Paper’s Organization
-Section 2 introduces the necessary background information for our domain of Wi-Fi–enabled battery-free systems.
-Our main problems are stated in Section 3, while Section 4 gives insight into how we solved these problems.
-Section 5 presents our evaluations on a widely used embedded platform (i.e., ESP32-C3).
-Related work is discussed in Section 6.
-Section 7 outlines future work and concludes the paper.
+The ESP32-Open-MAC project (Devreker, 2023) provides reverse-engineered Wi-Fi drivers for an Xtensa-based SoC, which we leveraged for the RISC-V–based ESP32-C3.
+Our work presents open-source Wi-Fi drivers for the ESP32-C3 alongside their static worst-case analyzability.
+In summary, our contributions are threefold:
+Case Study on Reverse Engineering: We present our reverse engineering efforts for the Wi-Fi driver of the ESP32-C3.
+Energy-Consumption & Execution-Time Model for Worst-Case Analysis: Our static analysis comes with a fine-grained energy-consumption model.
+Open-Source Implementation & Evaluation: Using the reverse-engineered knowledge, we develop an open-source Wi-Fi driver for the ESP32-C3 and present evaluation results that validate the driver’s property of being statically analyzable.
 2. Background
-The following section first gives some background regarding embedded network stacks (see Section 2.1).
-In section 2.2, we discuss mechanisms through which processors and Wi-Fi peripherals communicate in low-power SoCs.
-Finally, 2.3 describes the approach of worst-case resource-consumption analysis and its role in intermittent systems.
+This section gives background on network stacks and mechanisms for processor and Wi-Fi module communication. Finally, we outline worst-case resource-consumption analyses of program code.
 2.1. Network Stack
@@
 On top of the physical layer is the medium access control (MAC) that coordinates transmission between multiple devices.
-The MAC layer is tasked with network association, collision avoidance, or data encapsulation.
-In the case of Wi-Fi, collision avoidance (CSMA/CA) mechanism is used to prevent multiple devices from transmitting at the same time.
-Each layer of the network stack encapsulates data from the upper layer based on network configuration and passes it to the layer below.
-On intermittent devices, power may be lost at any point before packet transmission is completed, leading to loss of progress up to that point.
-In the case of Wi-Fi, PHY operations are implemented in hardware due to strict timing requirements.
-Wi-Fi peripherals that use the FullMAC architecture may implement both MAC and PHY layer operations in hardware.
-In some peripherals, tasks such as network association may be implemented in software (SoftMAC).
-Popularly, Wi-Fi devices operate in the basic service set (BSS) topology and connect to an access point (AP) that arbitrates communication within the network.
-Once connected, the devices are known as stations.
+In the case of Wi-Fi, the carrier sense multiple access/collision avoidance (CSMA/CA) mechanism is used to coordinate transmissions.
+Each layer of the network stack encapsulates data from the upper layer and passes it to the layer below.
+On intermittent devices, power may be lost at any point before packet transmission is completed, leading to a loss of progress up to that point.
+In the case of Wi-Fi, PHY-layer operations are implemented in hardware due to strict timing requirements.
+Some Wi-Fi peripherals implement MAC operations in software (SoftMAC), while others implement both MAC and PHY operations in hardware (FullMAC).
 2.2. Wi-Fi Hardware
@@
 Espressif devices such as ESP32-C3 follow this architecture.
-The peripheral communicates events such as successful transmission or reception of a packet through interrupts.
-On the other hand, external Wi-Fi chipsets may use the SPI interface for communication with the host controller.
-Besides the ESP32-C3’s Wi-Fi chip, the CYW43439 is another example suitable for low-power systems (Cypress Semiconductor Corporation, 2021).
-2.3. Static Worst-Case Analysis
-Static worst-case resource-consumption analysis comprises two major pillars: (1) a hardware-agnostic path analysis and (2) a hardware-dependent cost modeling capturing the actual resource demand (i.e., execution time or energy demand) on the target platform.
+The peripheral communicates events such as the successful transmission or reception of a packet through interrupts.
+On other platforms, external Wi-Fi chipsets may use the SPI interface to communicate with the host controller.
+2.3. Static Worst-Case Resource Analysis
+Static worst-case resource-consumption analysis of program code comprises two major pillars: (1) a hardware-agnostic program path analysis and (2) a hardware-dependent cost modeling capturing the actual resource demand (i.e., execution time or energy demand) for the actual target platform.
 The main idea is to formulate a mathematical problem formulation, usually an integer linear program (ILP): The ILP’s constraints originate from the path analysis.
+Abstract-interpretation techniques are further useful for refining the path constraints in a context-sensitive manner.
 The ILP’s objective function states the maximization of the actual costs with the given path constraints.
-Eventually, the ILP’s solution yields a mathematically proven worst-case bound, which, in turn, is necessary to ensure runtime guarantees.
+Eventually, the ILP’s solution yields a mathematically proven worst-case bound.
+With regard to energy consumption, the WCEC bound is useful to guarantee the execution without power failures during the runtime of batteryless devices.
 3. Problem Statement
@@
 3.1. Problem of Closed-Source Wi-Fi Drivers
-Our approach to intermittently-powered devices exploits static worst-case analysis techniques to gain knowledge on the worst-case energy consumption of operations.
-During runtime, our system’s scheduler checks the available energy and dispatches tasks only if sufficient energy is available to guarantee freedom from power failures.
-This WCEC-aware approach gives both memory-consistency and forward-progress guarantees.
-The trade-off we exploit is favoring expensive design-time analyses over complicated runtime checkpointing mechanisms.
+In intermittently-powered devices, static worst-case analysis techniques can be crucial to gain knowledge on the worst-case energy consumption of operations (Choi et al., 2022; Raffeck et al., 2024; Yarahmadi and Rohou, 2020).
+During runtime, the system’s scheduler checks the available energy (comparable to just-in-time checkpointing) and reactively dispatches tasks only if sufficient energy is available to guarantee freedom from power failures.
+The WCEC-aware approaches give both memory-consistency and forward-progress guarantees.
+The trade-off we exploit favors expensive design-time analyses over complicated runtime checkpointing mechanisms.
 Once deployed, systems achieve provable resilience against power failures.
 However, these design-time analyses require the code to be available.
-Our target platform is the ESP32-C3, a RISC-V–based, widely used, and low-cost SoC.
-Although BLE, likewise available on the ESP32-C3, is less power-intensive compared to Wi-Fi, we are interested in the more challenging Wi-Fi stack being of high relevance: In 2024, an estimated 31 % of IoT devices used Wi-Fi for connectivity (Sinha, 2024).
-For this ESP32-C3 platform, the Wi-Fi drivers were not available as analyzable source code, which is a common problem for embedded SoC platforms.
-Of particular relevance for our analysis are the MAC and PHY layer operations.
-Our Approach: To solve the problem of the closed source, our approach exploits several reverse-engineering techniques on Espressif’s closed-source Wi-Fi driver.
+Our target platform, the ESP32-C3, is a RISC-V–based, widely used SoC from Espressif, which has closed-source Wi-Fi drivers.
+Although less power-intensive networking technology exists (e.g., BLE), we are interested in the Wi-Fi stack being of high relevance: In 2024, an estimated 31 % of IoT devices used Wi-Fi for connectivity (Sinha, 2024).
+Further, the recent article on The Internet of Batteryless Things (Ahmed et al., 2024) calls for action and states that popular network protocols, including Wi-Fi, should be adapted for intermittent operation.
+Unfortunately, for our target, the generic Wi-Fi drivers were not available as analyzable source code, which is a common problem in embedded SoCs.
+Of main interest for our analysis are the MAC and PHY layers.
+Our Approach in a Nutshell: To solve the problem of the closed source, we exploit several reverse-engineering techniques on Espressif’s closed-source Wi-Fi driver alongside utilizing existing knowledge (Devreker, 2023).
 Fortunately, Espressif legally allows reverse engineering this driver.
@@
 3.2. Missing Resource-Consumption Model
-Although the developed open-source driver gives us the freedom to perform static analysis,
-a major problem is connecting memory-mapped IO operations to semantics with regard to the power demand.
-When considering the following code for modifying memory
-*((uint32_t*)0x60033ca0) &= 0xff00efff
-the static analysis is not aware of the semantics of this memory access.
-On the ESP32-C3 target platform, this operation switches on the transceiver from a sleep state to standby mode (see also Figure 1).
-As a consequence thereof, our static analysis of this memory access requires the notion that the power drawn increases after this operation.
-Similarly, specific instructions also require a longer execution time (than a normal RAM access), which in turn leads to higher energy demand (i.e., power over time).
-Consequently, we require a model that comprehensively accounts for power, energy, and time.
-Our Approach: For our target, we present a resource-consumption model that considers transitions in the hardware that change the power demand of the entire SoC.
+Although the developed open-source driver gives us the freedom to perform static analysis, a major problem is connecting memory-mapped IO operations to semantics with regard to the power demand.
+When considering the following MMIO operation:
+*((uint32_t*)0x60033ca0) &= 0xff00efff,
+the static analysis is not aware of its semantics. On the ESP32-C3, this operation switches the transceiver from a sleep state to standby mode (see also Figure 2).
+Hence, our static resource analysis needs to include the notion that the power drawn increases after this operation.
+Similarly, specific instructions also require a longer execution time than a normal RAM access, which in turn leads to higher energy demand (i.e., power over time).
+Consequently, a comprehensive model needs to account for power, energy, and time.
+Our Approach in a Nutshell: For our target, we present a resource-consumption model that considers transitions in the hardware, which change the SoC’s power demand.
 Our analysis makes use of this model and tracks program paths in a context-sensitive way.
 4. Approach
-We use a combination of commonly available static and dynamic reverse-engineering analysis methods for identifying the registers and correspondingly their contents.
-Espressif provides their SoftMAC implementation as binaries under the permissive Apache license.
-The binaries also contain the symbol table.
+We use a combination of commonly available static and dynamic reverse-engineering analysis methods for identifying the registers and, correspondingly, their contents.
 4.1. Reverse Engineering
 Static Analysis
-Through decompilation, the execution flow and accordingly the registers used for transmission and reception can be identified.
+Through decompilation, the program flow and, accordingly, the registers used for transmission and reception can be identified.
 Functions within the binaries are named based on their purpose, which leads to informative decompilation results.
-Additionally, the interrupt configuration used by the Wi-Fi peripheral can be identified as this is not exposed to developers.
+Additionally, the undocumented interrupt configuration used by the Wi-Fi peripheral can be identified.
 However, decompilation alone is insufficient to identify control and data flows.
 Several of the registers are configured based on data that is internal to the driver, whose structure is not documented.
-Moreover, the peripheral also has several slots for transmission, which is also dynamically selected based on the network configuration.
+Moreover, the peripheral also has several slots for transmission, which are dynamically selected based on the network configuration.
 Therefore, we use QEMU to dynamically determine the register contents as well as the slots.
 Dynamic Analysis
-The QEMU version in (lcgamboa, 2023), a fork of Espressif’s QEMU, supports Wi-Fi for ESP32-C3 by identifying a select set of registers and emulating an open Wi-Fi AP.
-We modify this fork to log all peripheral accesses and their contents.
-This allows us to verify register addresses as well as determine the register operations corresponding to the Wi-Fi network configuration in QEMU.
-While QEMU provides insights into all the peripheral’s register accesses and their contents, its emulation does not fully account for peripheral configuration.
-It only emulates the DMA transfer as well as the interrupt mechanism.
+The QEMU version in (lcgamboa, 2023), a fork of Espressif’s QEMU, supports Wi-Fi for ESP32-C3 by emulating an open Wi-Fi access point (AP).
+In particular, it only implements registers required for triggering a transmission and reading received packets through DMA as well as the status register, which encodes the reasons for an interrupt occurrence.
+We further modify this fork to log all register accesses and their contents within the Wi-Fi peripheral’s address range.
+While the QEMU fork provides insights into accesses of the peripheral’s registers and their contents used during the emulation, it does not fully emulate the hardware.
+Hardware initialization, for example, is not emulated.
 Moreover, it is not always possible to deduce the purpose of the register operation.
-Binary Patching for Injecting Tracing Code
-We modify the control flow to trace execution in real-time.
-We make no modifications to the binaries themselves but replace exported symbol names within some binary files.
-The new symbol name can then be declared as a function as part of a regular ESP-IDF sketch and can call the original function with the provided arguments.
-This also allows us to peek into the arguments themselves without using breakpoints or external debuggers.
-The resulting binary can also be run on the emulator or on the ESP32-C3.
-However, this only allows us to create such function redirects only for those functions that are declared in one object file but called from another.
-These reverse-engineering methods in combination with the existing infrastructure in ESP32-Open-MAC yield necessary information for constructing the Wi-Fi driver for ESP32-C3.
-[device graph]device graph showing energy states
+Binary Hooking for Injecting Tracing Code
+The modules of Espressif’s Wi-Fi driver are provided as archives (.a) of object (.o) files.
+These object files have undefined symbols referring to functions defined in other modules, such as the binary blobs of the ESP32-C3 ROMs.
+Their definition is resolved during the final linking of the executable.
+To trace the execution, we hook into these symbols by replacing the symbol names and exploiting the original symbol names in our custom firmware to identify the control flow.
+These reverse-engineering methods, in combination with the existing infrastructure in ESP32-Open-MAC, yield the necessary information for constructing the Wi-Fi driver for the ESP32-C3.
 4.2. The Open-Source ESP32-C3 Driver
-Once the registers and transmit/receive mechanisms are identified, we use the software architecture shown in Figure 2 to develop the driver.
+Once the registers and transmit/receive mechanisms are identified, we use the architecture shown in Figure 1 to develop the driver.
 MAC Task
@@
 Driver Task
-The driver task interacts with the MMIO registers.
-It also manages the memory allocated to DMA structures required by the peripheral.
-All memory is statically allocated to aid WCET/C analysis.
+The driver task interacts with the MMIO registers and manages memory allocated to DMA structures required by the peripheral.
+All memory is statically allocated to aid WCET/WCEC analysis.
 Processing of interrupts is deferred by the interrupt handler to the driver task.
-Incoming 802.11 packets are passed to the MAC layer for further processing.
-The driver task also defines two queues: a hardware event queue and a MAC event queue.
-The interrupt handler notifies the driver of the occurrence and reason for the interrupt using the hardware event queue.
+Incoming 802.11 packets are passed to the MAC task for further processing through a queue.
+The interrupt handler notifies the driver of the occurrence and reason for the interrupt using a hardware event queue.
 The MAC task also writes packets to be transmitted into this queue.
@@
 4.3. Building a Resource-Consumption Model
-For tasks such as transmission and reception, the Wi-Fi peripheral functions independently of the CPU.
-Therefore, static code analysis has to be supplemented with resource consumption models, namely time and energy.
+For device operations such as transmission and reception, the Wi-Fi peripheral operates independently of the CPU.
+To yield sound results, static resource analysis requires a cost model specifying the maximum time and energy demands of operations.
+These paragraphs provide a generic overview of the cost model, while Section 5.2 discusses concrete values.
 Timing Information
-The Wi-Fi peripheral performs packet transmission and notifies the CPU of the status through interrupts.
-Similarly, some MMIO register operations take time due to ongoing receive tasks.
-However, these tasks are still completed or fail within a finite amount of time and therefore have upper bounds on their execution times.
+The necessary timing-related costs comprise both protocol- and hardware-specific aspects.
+The application-specific protocol influences the maximum duration of device operations (e.g., through the maximum packet size).
+Additionally, timeouts and operation windows in the protocol allow deriving an upper bound on operations’ execution times, such as waiting to receive a packet.
+Bounds on the duration of operations also enable bounding the occurrence of interrupts related to the completion of a transmission or an incoming packet.
+The second aspect, the hardware-specific behavior, concerns the Wi-Fi peripheral’s MMIO interface.
+The execution time of certain MMIO operations, such as waiting for a bit to change its value, which signals the completion of a configuration update, depends on the reaction of the hardware.
+[device graph]device graph showing energy states
 Energy Information
-The device energy model should consist of all possible energy states, as well as the corresponding energy consumption.
-In our current setup, we assume that the CPU is in the always-on state; however, modeling the ESP32-C3’s sleep states, (low-power) run states, and frequency changes is possible (Dengler et al., 2023).
-For the Wi-Fi peripheral, we define three states of operation: Sleep, Standby and Transmitting.
-The Wi-Fi peripheral is powered-down in and therefore consumes no energy.
-In , the Wi-Fi peripheral is powered up and may receive packets, thereby consuming energy.
-In , the device transmits packets.
-The energy consumption varies based on the duration of transmission.
-No reception occurs in this state, as the peripheral owns the channel through the CSMA-CA mechanism.
-The device state machine is shown in Figure 1.
+The energy-related cost model for the Wi-Fi peripheral consists of all possible device states with distinct power consumption.
+A simplified version of the device graph modeling the states and their transitions for the Wi-Fi peripheral is shown in Figure 2.
+The depicted graph features three states of operation: Sleep, Standby, and Transmitting.
+The Wi-Fi peripheral is powered down and inoperative in the sleep state ( ).
+In the standby state ( ), the Wi-Fi peripheral is powered up and may receive packets.
+Lastly, in the transmitting state ( ), the device actively transmits a packet.
+No reception occurs in this state, as the peripheral owns the channel through the CSMA/CA mechanism.
+For this paper, we assume that the CPU is always on and uses only one run mode, although modeling the ESP32-C3’s sleep states, run modes, and clock-frequency changes is generally possible (Dengler et al., 2023).
 4.4. Implementation
 Reverse Engineering
-The ESP32-Open-MAC project lists the registers that are necessary to achieve basic transmission and reception on the ESP32.
-In order to identify the equivalent registers on the ESP32-C3, we build a sketch that performs simple UDP transmission and analyze it using Ghidra’s (Agency, 2019) decompiler.
-We also add volatile memory sections for the peripheral ranges identified in previous reverse engineering efforts (lcgamboa, 2023).
-The author also provides a fork of Espressif’s QEMU that adds support for Wi-Fi functionality on the ESP32-C3.
-We modify this fork to log all peripheral accesses.
-We replace only the external function names in the binaries with function names of similar length.
-Furthermore, we declare these functions in the main ESP-IDF sketch to the original function names in the object files.
-The resulting sketch is both loaded to the emulator and flashed to an ESP32-C3-Devkit-M1.
+The ESP32-Open-MAC project lists the necessary MMIO registers to achieve basic transmission and reception on the ESP32.
+On the ESP32, the MMIO addresses belong to the 0x3ff73000-0x3ff75fff range, but on the ESP32-C3 are mapped to 0x60033000-0x60035fff.
+The offsets of ESP32-C3’s MMIO registers from the base address differ from those of the ESP32.
+In order to identify the equivalent registers on the ESP32-C3, we write a firmware that performs simple UDP transmission and analyze it using Ghidra’s (Agency, 2019) decompiler.
+In Ghidra, we mark the memory ranges of the peripheral identified with QEMU (lcgamboa, 2023) as volatile and identify the functions that access these memory locations.
+Functions with prefixes such as hal_mac_* or wdev_mac_* were found to handle MAC/PHY layer operations, which are of particular interest for our reverse engineering.
+Among the identified functions, we hook into those that are listed as undefined symbols in object files within the libpp.a archive of the proprietary blobs.
+The resulting firmware with the hooks is loaded into our QEMU fork.
 | Register | Deduced Purpose | Access |
@@
 | 0x60033cac | Clear Tx slot | R/W |
-Some important registers and their deduced purpose are shown in Table 1.
+An excerpt of the most important registers identified through our reverse engineering is shown in Table 1, along with the deduced purpose of these registers.
 Interrupt Handling
-Setting up interrupts on the ESP32-C3 is well documented.
-However, the Wi-Fi interrupt is not exposed to developers.
-Through decompilation, we find that the Wi-Fi driver uses CPU interrupt number 1.
-To register a custom Wi-Fi interrupt handler, we find that first the source register corresponding to interrupt 1 (0x600c2000) has to be cleared.
-The interrupt has to be disabled by masking bit 1 of 0x600c2104.
-The custom interrupt handler can be set using intr_handler_set provided by ESP-IDF.
-Finally, the interrupt is enabled again by setting bit 1 of 0x600c2104.
+The details of the Wi-Fi interrupt handling are not exposed to developers.
+Through decompilation, we find that the Wi-Fi driver uses CPU interrupt number 1, which is marked as reserved in the manual (Espressif Systems, 2025b), and uses an undocumented register for that interrupt source at address 0x600c2000.
+With this information, registering a Wi-Fi interrupt handler follows the usual procedure on the ESP32-C3.
+First, the (undocumented) interrupt source register (0x600c2000) is cleared, and the interrupt is disabled by clearing bit 1 of the (documented) interrupt-enable register (0x600c2104).
+The interrupt handler is registered using ESP-IDF’s intr_handler_set.
+Finally, the interrupt is enabled by setting bit 1 of 0x600c2104 and loading the interrupt number to 0x600c2000.
 ESP32-C3 Wi-Fi Driver
-We implement the driver using the architecture described in Figure 2.
-We use an open AP, as the encryption functionality of Wi-Fi peripheral hardware is not fully reverse-engineered.
-For initialization and connecting to an AP, we use the API calls esp_wifi_init and esp_wifi_start provided by ESP-IDF.
-These calls initialize and configure the Wi-Fi peripheral, set the internal interrupt handler for Wi-Fi, and initiate a connection to the AP.
-The successful completion of these tasks can be monitored through a system notification WIFI_EVENT_STA_CONNECTED.
-Once the notification is received, we kill the Wi-Fi task using the undocumented pp_post function and set up the custom interrupt handler.
+We implement the driver using the architecture shown in Figure 1.
+We use an open AP, as the encryption functionality of the Wi-Fi hardware is not fully reverse-engineered.
+For initialization and connecting to an AP, we use ESP-IDF’s API.
+The functions esp_wifi_init and esp_wifi_start initialize and configure the Wi-Fi, set the internal interrupt handler for Wi-Fi, and initiate an AP connection.
+Their successful completion can be monitored through the system notification WIFI_EVENT_STA_CONNECTED.
+Once the notification is received, the Wi-Fi task can be killed using the undocumented pp_post function, which queues instruction codes to the Wi-Fi task started by the driver.
 Finally, our driver and MAC tasks are started.
 We use PfIP (Vogelgesang et al., 2025) as the IP stack due to its intermittency-aware execution.
-Static Analysis
-For static analysis, we use the Platin framework (Maroun et al., 2024).
-Platin uses a modified version of the clang compiler that collects program flow information.
-The compiler generates a metadata file from the program flow information and transforms it into an ILP, which, in turn, can be solved using programs such as lp_solve (Berkelaar et al., 2004) or gurobi (Gurobi Optimization, LLC, 2023).
-For the purpose of this paper, we annotate the current consumption for each state in the device’s energy model based on the datasheet (Espressif Systems, 2022).
-Additionally, Platin distinguishes between the energy penalty of switching on and off any given energy state (wcec-turn-on and wcec-turn-off) and the energy consumption of a device in the state energy-stay-on.
-wcec-always-on represents the energy required when the device is in any given state.
-In our model, we set both the switching on and off penalties to zero and model the device-switching behavior pessimistically.
 5. Evaluation
-This section presents the evaluation of our open-source Wi-Fi driver.
+We perform a static resource analysis of our Wi-Fi driver to assess the ability to bound the program flow and resource demand.
 Section 5.1 describes the hardware and network setup as well as the task system used for the experiments.
 Additionally, we detail the required assumptions for the static resource-consumption analysis using the Platin analysis framework (Maroun et al., 2024).
-In Section 5.2, we then examine and discuss the results from the static analysis.
+In Section 5.2, we examine and discuss the analysis results.
 5.1. Evaluation Setup
@@
 Functionality
-To validate the functionality of the Wi-Fi driver, we perform an experiment with three tasks.
-The mac_task interacts with the PfIP network stack, the hw_task initiates the transmission of packets required by the mac_task and handles receiving, and a test_task sends out a packet using the PfIP stack and waits for an answer from other devices in the network.
-We validated the functionality of our open-source driver with Wireshark captures of the network traffic.
-Static Analysis
-To analyze the worst-case resource demand of our Wi-Fi driver, we use the Platin analysis framework (Maroun et al., 2024) in combination with the device-model analysis of the WoCA approach (Raffeck et al., 2024).
-The focus of our evaluation lies solely on the driver functions, independent of the runtime and operating system.
-While a whole-system analysis, which includes the operating system, is generally possible (Schuster et al., 2019), we exclude the FreeRTOS components and the interaction with the FreeRTOS scheduler from our analysis.
-Likewise, we exclude the callbacks into the PfIP network stack from the analysis.
-We perform the worst-case resource-demand analysis both for individual driver functions and for a task comprising multiple device interactions.
-For each individual driver function, we determine an upper bound on the execution time.
-The evaluation task transmits a packet and actively waits for the completion of the transmission.
+To validate the driver’s functionality, we write a firmware that sends UDP packets through PfIP’s network stack.
+PfIP is provided with the mac_task’s transmit handle, which in turn queues the packet for physical transmission by the hw_task.
+The responses from the network are passed back through PfIP’s receive callback.
+We capture the network traffic using Wireshark to monitor our driver’s functional correctness.
+Static Resource Analysis
+To analyze the worst-case resource demand of our Wi-Fi driver, we use the Platin analysis framework (Maroun et al., 2024).
+With the device-model analysis of the WoCA approach (Raffeck et al., 2024), Platin supports device-aware resource-consumption analysis for the ESP32-C3.
+Platin uses a modified version of the clang compiler that collects user-annotated knowledge, derives metainformation about the program flow, and tracks the relation between source and object level regarding this control-flow information.
+Using this metainformation and hardware-specific cost models, Platin formulates the costs of and constraints on the program behavior as an ILP suitable for common mathematical solvers, such as lp_solve (Berkelaar et al., 2004) or Gurobi (Gurobi Optimization, LLC, 2025), to obtain the maximum resource demand.
+For the purpose of this paper, we annotate the current consumption for each state in the device’s energy model based on the values provided by the manufacturer in the datasheet (Espressif Systems, 2022).
+For the processor, we rely on the model for the ESP32-C3 already present in Platin.
+The focus of our evaluation lies on the driver functions, independent of the runtime and operating system.
+While a whole-system analysis, which includes the operating system, is generally possible (Schuster et al., 2019), we exclude the FreeRTOS components and callbacks into the PfIP network stack from our analysis for this paper.
+We analyze benchmarks including individual driver functions and an evaluation task comprising multiple device interactions.
+For each driver function, we determine an upper bound on the execution time.
+The evaluation task sends a packet and actively waits for the transmission’s completion.
 After starting the transmission, the Wi-Fi transceiver transmits the packet and waits for an acknowledgment (ACK).
@@
 […]
-Most functions of our open-source Wi-Fi driver perform only operations on memory-mapped registers to control the transceiver and exhibit relatively low execution time and energy consumption.
-wifi_mac_handle_rx copies the payload data of the received package between buffers, so the analysis has to assume the worst case of the maximum packet size.
-Similarly, the analysis has to assume the worst-case number of packets to handle for the execution time of wifi_handle_rx and the maximum time until an ACK is received for the execution time of wifi_wait_for_tx.
-Finally, the difference between the always-on and device-aware WCEC bounds underlines the importance of considering the states of the Wi-Fi transceiver in a fine-granular manner.
+Most functions of our open-source Wi-Fi driver perform only operations on memory-mapped registers to control the transceiver and exhibit a relatively low execution time and energy consumption.
+The few notably higher resource bounds relate to worst-case analysis assumptions for the control flow.
+wifi_mac_handle_rx moves the payload of the received package between buffers, so the analysis has to assume the worst case of the maximum packet size.
+Similarly, the analysis has to assume the worst-case number of packets for the execution time of wifi_handle_rx and the maximum time until an ACK is received for wifi_wait_for_tx.
+The difference between the always-on and device-aware WCEC bound underlines the importance of considering the states of the Wi-Fi transceiver in a fine-granular manner.
+We demonstrated that the reverse engineering of our driver code enables deriving the necessary control-flow bounds and performing accurate static resource analysis of Wi-Fi operations on the ESP32-C3.
 6. Related Work
-Typical approaches for battery-free wireless communication use Bluetooth Low Energy (BLE), LoRa , or backscatter-based wireless communications (Cai et al., 2023; Talla et al., 2017; Jiang et al., 2023).
-Intermittency-aware operation has been successfully demonstrated, especially for BLE and LoRa.
 Intermittency-Aware Wireless Communication
-Wireless networking is a key challenge in Battery-free IoT designs due to intermittency.
-Therefore, the networking stack deployed on these devices must be capable of intermittent execution.
-(Fu et al., 2023) proposed a fully intermittency-aware networking stack.
-Schoeberl et al.
-developed tpIP, a time-predictable IP stack for cyber-physical networks (Schoeberl and Pedersen, 2018).
-tpIP is designed to allow static analysis through bounded loops and avoids dynamic memory allocation.
-They also estimated WCET for the stack’s operations.
+Wireless networking is a key challenge in batteryless designs due to intermittency.
+Networking stacks deployed on such devices must be fully aware of intermittent execution (Fu et al., 2023).
+In contrast to our work on Wi-Fi, other approaches for batteryless wireless communication use Bluetooth Low Energy (BLE), LoRa, or backscatter-based wireless communications (Cai et al., 2023; Talla et al., 2017; Jiang et al., 2023).
+As a time-predictable IP stack, tpIP (Schoeberl and Pedersen, 2018) is designed to allow static analysis through bounded loops and avoids dynamic memory allocation.
 Similarly, Vogelgesang et al.
-designed PfIP, a transactional networking stack with worst-case bounds for every transaction (Vogelgesang et al., 2025).
+designed PfIP, a networking stack with transactional semantics and worst-case bounds for every transaction (Vogelgesang et al., 2025).
+Transactional semantics break down the overall task into atomic sub-tasks (or transactions) that can be independent of each other.
 These approaches are agnostic to the underlying physical layer.
-De Winkel et al.
-introduced a battery-free smartwatch running FreeBie (de Winkel et al., 2022), a bidirectional BLE stack for intermittently-powered Bluetooth devices.
-The device stored checkpoints for all processes when power failure was imminent and restored them when power was available.
-Additionally, due to intermittency, two devices may not be awake at the same time for communication.
-In (Geissdoerfer and Zimmerling, 2021, 2022), these challenges are addressed through probabilistic modeling of latency and charge times, respectively.
-Similarly, a LoRa-based battery-free image sensing system that adaptively configures its transmission mode based on the available energy was developed in (Nardello et al., 2019).
-Raffeck et al.
-demonstrated a LoRa-based device using WoCA, a whole system worst-case analysis approach for intermittent wireless devices (Raffeck et al., 2024).
-For static analysis, Platin was used with loop bounds calculated based on the LoRa device configuration.
-Greentooth is a proposed low-power wireless protocol designed for battery-free devices (Babatunde et al., 2024).
-Greentooth synchronizes communication between two devices using time division multiple access (TDMA)-style scheduling and low-power wake-up radios.
-Wi-Fi remains under-explored due to its higher power requirement compared to BLE or LoRa.
-The ubiquity of Wi-Fi networking applications and availability of off-the-shelf inexpensive hardware such as the ESP32-C3 promises battery-free innovation in the field of, for example, smart-home applications, where an existing Wi-Fi network is usually available.
+FreeBie (de Winkel et al., 2022) is a bidirectional BLE stack designed for intermittent communication, demonstrated with a batteryless smartwatch.
+Cameroptera (Nardello et al., 2019) is a LoRa-based batteryless image sensing system that adaptively configures its transmission mode based on the available energy.
+RockClimb (Choi et al., 2022) uses WCET bounds without device awareness to guarantee the reaching of checkpoints.
+Yarahmadi et al.
+(Yarahmadi and Rohou, 2020) employ WCEC-aware checkpoint placement without considering the resource demand of peripheral devices.
+WoCA introduces device-aware whole-system worst-case analysis for intermittent systems, including a model for LoRa-based communication (Raffeck et al., 2024).
+Without an intermittency-aware protocol, two devices may not be awake at the same time for communication.
+Greentooth (Babatunde et al., 2024) synchronizes communication between two devices using time division multiple access (TDMA)-style scheduling and low-power wake-up radios.
+Probabilistic modeling of latency or charge times (Geissdoerfer and Zimmerling, 2021, 2022) likewise achieves synchronized communication.
+For intermittent devices, Wi-Fi is under-explored due to its high power demand but promises batteryless innovation, for example, in fields like smart-home applications.
 Open-Source Wi-Fi Platforms
-Open-source drivers exist for several Wi-Fi network interface cards from, for example, Atheros or Mediatek (Project, 2024).
-These require no binary blobs and are often supported by the manufacturers themselves.
-However, they require a Linux kernel and PCI-E hardware support and are generally not intended for low-power SoCs.
+Outside the world of low-power SoCs, open-source drivers exist for several Wi-Fi network interface cards, often supported by the manufacturers themselves, for example, from Atheros or Mediatek (Project, 2024).
 The OpenWiFi project (Jiao et al., 2020) developed an open-source FPGA-based software-defined radio for the 802.11 standard paired with an RF frontend.
-However, FPGAs are not commonly used in consumer hardware and the platform is intended to be a tool for researchers.
-On low-power devices, the ZeroWi project (Bentham, 2020) focuses on reverse-engineering the capabilities of the Cypress CYW43438 Wi-Fi SoC.
-The CYW43438 consists of an internal ARM Cortex CPU, which is programmed using non-free firmware provided by Cypress.
+On low-power devices, the ZeroWi project (Bentham, 2020) reverse-engineers the capabilities of the Cypress CYW43438 Wi-Fi SoC, which is programmed using non-free firmware.
 Drivers written for this chip then send commands in the SDIO format over SPI to communicate with the firmware.
-For the Espressif products, early attempts at reverse engineering were performed by Uri Shaked, who used it as part of his Wokwi simulator for ESP32 (Shaked, 2019).
-The QEMU-based open-source PICSimLab simulator introduces support for Wi-Fi emulation on the ESP32 and ESP32-C3 devices (lcgamboa, 2022).
+Early attempts at reverse engineering Espressif’s Wi-Fi driver were performed by Uri Shaked as part of the Wokwi simulator for ESP32 (Shaked, 2019).
+The QEMU-based open-source PICSimLab simulator introduces support for Wi-Fi emulation on the ESP32 and ESP32-C3 (lcgamboa, 2022).
 Based on these attempts, the ESP32-Open-MAC project (Devreker, 2023) aims to fully open-source the Wi-Fi driver on the ESP32.
-Our work not only ported ESP32-Open-MAC onto the RISC-V–based ESP32-C3 but also integrated it with (1) a resource-consumption model for static analysis and (2) a transactional network stack that is suitable for battery-free devices.
+Our work not only ported ESP32-Open-MAC onto the RISC-V–based ESP32-C3 but also integrated it with a resource-consumption model for static analysis and a transactional network stack that is suitable for batteryless devices.
 7. Future Work & Conclusion
-The current energy model does not consider packet reception.
-Therefore, future work needs to consider a more detailed device model: That is, pessimism in the bounds can be reduced by a more fine-grained device model.
-For example, in , the peripheral is not actively transmitting after the end of transmission until ACK is received.
+The use of Wi-Fi for intermittently-powered batteryless systems is an ongoing research challenge since network protocols are not designed with intermittent operation in mind (Ahmed et al., 2024).
+With this work, we contribute to this challenge with our reverse-engineered and publicly-available Wi-Fi drivers, along with their static worst-case energy-consumption analysis.
+Our analysis results allow intermittent systems to execute with runtime guarantees on the power-failure–free execution of operations.
+Future work needs to consider a more detailed device model to reduce the pessimism in the bounds.
+For example, in , the peripheral is not actively transmitting after the end of transmission until the ACK is received.
 Additionally, energy penalties for switching on/off energy states could be accurately considered.
-Further, we use a static peripheral configuration.
+We also use a static peripheral configuration.
 The resource-consumption bounds are impacted by peripheral configurations, such as data rate of transmission.
-An interesting avenue for future work is to identify optimal configurations under energy constraints.
-The ESP32-Open-MAC project also has recently introduced initialization independent of FreeRTOS.
+Future work could identify optimal configurations under energy constraints.
+ESP32-Open-MAC has recently introduced an initialization of the Wi-Fi peripheral independent of FreeRTOS.
 Therefore, an OS-aware implementation could also result in easier static analysis, for example, by exploiting Rust’s statically analyzable async/await pattern.
 In this work, we presented a fully open-source Wi-Fi driver for the ESP32-C3, based on the ESP32-Open-MAC project.
-We successfully port the project from the Tensilica-based ESP32 to the RISC-V–based ESP32-C3 using common reverse-engineering techniques.
-Finally, we present a proof-of-concept for performing static resource-consumption analysis on the physical operations of the Wi-Fi peripheral using the extended Platin worst-case analysis toolkit.
-For the future, we envision more exchange between the static-analysis and the security communities.
-More interdisciplinary research on this topic could avoid a future of insecure and unsafe battery-free IoT systems.
-[ backgroundcolor=green!10, linecolor=green!70,linewidth=2pt,topline=false,bottomline=false, innertopmargin=0pt, innerbottommargin=5pt, innerleftmargin=6pt, innerrightmargin=6pt, frametitleaboveskip=5pt, frametitlefont=, frametitlealignment=, frametitle=The implementation is available at: https://git.nt.uni-saarland.de/open-access/pfip ]
+We successfully port the project from the Xtensa-based ESP32 to the RISC-V–based ESP32-C3 using common reverse-engineering techniques.
+Finally, we present a proof-of-concept for performing static resource-consumption analysis on the Wi-Fi peripheral’s physical operations using the extended Platin worst-case analysis toolkit.
+[ backgroundcolor=tolerableGreen!25, linecolor=tolerableGreen,linewidth=2pt,topline=false,bottomline=false, innertopmargin=0pt, innerbottommargin=3pt, innerleftmargin=6pt, innerrightmargin=6pt, frametitleaboveskip=3pt, frametitlefont=, frametitlealignment=, frametitle=The implementation is available at: ] https://git.nt.uni-saarland.de/open-access/pfip
 Acknowledgements.
-This work is supported by the German Research Foundation (Deutsche Forschungsgemeinschaft, DFG) as part of SPP 2378 (Resilient Worlds: project number 502615015, Resilient Power-Constrained Embedded Communication Terminals, ResPECT).
+This work is supported by the German Research Foundation (DFG) as part of SPP 2378 (Resilient Worlds: project number 502615015, Resilient Power-Constrained Embedded Communication Terminals, ResPECT) and the NGI0 Core Fund under grant agreement No 101092990.
 We would also like to thank all ESP32-Open-MAC contributors for their valuable contributions to the community.References
@@
 In Proceedings of the 10th International Symposium on Industrial Embedded Systems (SIES ’15).
-IEEE, 39–48.
+39–48.
 https://doi.org/10.1109/SIES.2015.7185039
@@
 Greentooth: Robust and Energy Efficient Wireless Networking for Batteryless Devices.
-ACM Transactions on Sensor Networks 20, 3 (may 2024), 66:1–66:31.
+ACM Transactions on Sensor Networks 20, 3 (2024), 66:1–66:31.
 https://doi.org/10.1145/3649221
@@
 Battery-Free Wireless Sensor Networks: A Comprehensive Survey.
-IEEE Internet of Things Journal 10, 6, March 15 (mar 2023), 5543–5570.
+IEEE Internet of Things Journal 10 (2023), 5543–5570.
 https://doi.org/10.1109/JIOT.2022.3222386
@@
 In Proceedings of the 28th Real-Time and Embedded Technology and Applications Symposium (RTAS ’22).
-IEEE, 40–54.
+40–54.
 https://doi.org/10.1109/RTAS54340.2022.00012
-- Cypress Semiconductor Corporation (2021) Cypress Semiconductor Corporation 2021.
-CYW43439 Single-Chip IEEE 802.11 b/g/n MAC/Baseband/Radio with Integrated Bluetooth 5.2 Compliance.
-Cypress Semiconductor Corporation.
-https://www.mouser.com/datasheet/2/196/Infineon_CYW43439_DataSheet_v03_00_EN-3074791.pdf
 - de Winkel et al.
@@
 In Proceedings of the 20th Annual International Conference on Mobile Systems, Applications and Services (MobiSys ’22) (Portland, Oregon) (MobiSys ’22).
-ACM, New York, NY, USA, 287–301.
+287–301.
 https://doi.org/10.1145/3498361.3538934
@@
 6:1–6:24.
-https://drops.dagstuhl.de/opus/volltexte/2023/18035/pdf/LIPIcs-ECRTS-2023-6.pdf
+https://doi.org/10.4230/DARTS.9.1.2
 - Devreker (2023) Jasper Devreker. 2023. Unveiling secrets of the ESP32: creating an open-source MAC Layer. Retrieved 2025-01-03 from https://zeus.ugent.be/blog/23-24/open-source-esp32-wifi-mac/
-- Espressif Systems (2022) Espressif Systems 2022. ESP32-C3 Series Datasheet. Espressif Systems. https://www.espressif.com/sites/default/files/documentation/esp32-c3_datasheet_en.pdf
+- Espressif Systems (2022) Espressif Systems. 2022. ESP32-C3 Series Datasheet. https://www.espressif.com/sites/default/files/documentation/esp32-c3_datasheet_en.pdf
+- Espressif Systems (2025a) Espressif Systems. 2025a. ESP IoT Development Framework. https://www.espressif.com/en/products/sdks/esp-idf
+- Espressif Systems (2025b) Espressif Systems. 2025b. ESP32-C3 Technical Reference Manual. https://www.espressif.com/sites/default/files/documentation/esp32-c3_technical_reference_manual_en.pdf
 - Fu et al.
@@
 In 18th USENIX Symposium on Networked Systems Design and Implementation (NSDI 21).
-USENIX Association, 439–455.
+439–455.
 https://www.usenix.org/conference/nsdi21/presentation/geissdoerfer
@@
 In 19th USENIX Symposium on Networked Systems Design and Implementation (NSDI 22).
-USENIX Association, 419–435.
+419–435.
 https://www.usenix.org/conference/nsdi22/presentation/geissdoerfer
-- Gurobi Optimization, LLC (2023) Gurobi Optimization, LLC. 2023. Reference Manual. gurobi.com/.
+- Gurobi Optimization, LLC (2025) Gurobi Optimization, LLC. 2025. Reference Manual. gurobi.com
 - Jiang et al.
@@
 openwifi: a free and open-source IEEE802.11 SDR implementation on SoC.
-In 91st IEEE Vehicular Technology Conference, VTC Spring 2020, Antwerp, Belgium, May 25-28, 2020.
-IEEE, IEEE, 1–2.
+In Proceedings of the 91st IEEE Vehicular Technology Conference (VTC 2020).
+IEEE, 1–2.
 https://doi.org/10.1109/VTC2020-Spring48590.2020.9128614
@@
 121.
-Schloss Dagstuhl - Leibniz-Zentrum für Informatik, 2:1–2:14.
+2:1–2:14.
 https://doi.org/10.4230/OASIcs.WCET.2024.2
@@
 In Proceedings of the 17th European Workshop on Systems Security (EuroSec ’24).
-ACM, 29–36.
+29–36.
 https://doi.org/10.1145/3642974.3652283
@@
 In Proceedings of the 7th International Workshop on Energy Harvesting & Energy-Neutral Sensing Systems (EnsSys ’19).
-ACM, 8–14.
+8–14.
 https://doi.org/10.1145/3362053.3363491
@@
 In Proceedings of the 25th ACM SIGPLAN/SIGBED International Conference on Languages, Compilers, and Tools for Embedded Systems (LCTES ’24).
-ACM, 83–94.
+83–94.
 https://doi.org/10.1145/3652032.3657569
@@
 In 2018 IEEE 21st International Symposium on Real-Time Distributed Computing (ISORC).
-IEEE, IEEE Computer Society, 75–82.
+IEEE, 75–82.
 https://doi.org/10.1109/ISORC.2018.00018
@@
 Dissertation.
-Darmstadt University of Technology, Germany, Darmstadt.
+Darmstadt University of Technology, Germany.
 http://tuprints.ulb.tu-darmstadt.de/7243/
@@
 In Proceedings of the 25th Real-Time and Embedded Technology and Applications Symposium (RTAS ’19).
-IEEE, 318–330.
+318–330.
 https://doi.org/10.1109/RTAS.2019.00034
@@
 - Sinha (2024) Satyajit Sinha. 2024. State of IoT 2024: Number of Connected IoT Devices Growing 13% to 18.8 Billion Globally. https://iot-analytics.com/number-connected-iot-devices/.
-- Systems (2016) Espressif Systems. 2016. ESP IoT Development Framework | Espressif Systems. https://www.espressif.com/en/products/sdks/esp-idf.
 - Talla et al.
@@
 Battery-Free Cellphone.
-Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies 1, 2 (Jun 2017), 25:1–25:20.
+Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies 1, 2 (2017), 25:1–25:20.
 https://doi.org/10.1145/3090090
@@
 In Proceedings of the Consumer Communications & Networking Conference (CCNC 2025).
+https://www.nt.uni-saarland.de/publications/vogelgesang2025pfip/PfIP_authors_version.pdf
+- Wägemann et al.
+(2018) Peter Wägemann, Christian Dietrich, Tobias Distler, Peter Ulbrich, and Wolfgang Schröder-Preikschat.
+2018.
+Whole-System Worst-Case Energy-Consumption Analysis for Energy-Constrained Real-Time Systems.
+In Proceedings of the 30th Euromicro Conference on Real-Time Systems (ECRTS ’18), Vol.
+106.
+24:1–24:25.
+https://doi.org/10.4230/LIPIcs.ECRTS.2018.24
 - Wilhelm et al.
@@
 12471.
-Springer, 169–185.
+169–185.
 https://doi.org/10.1007/978-3-030-60939-9_12